# Heun-function analysis of the Dirac spinor spectrum in a sine–Gordon soliton background

### v1

@@ 3 Wronskian method and matching scattering states
The solutions to the differential equation (2.13) can be expressed in terms of the local Heun functions [4, 7, 5]. Note that we have used the parameter set provided in (2.52)-(2.54) for $e_{1}e_{2}=-1,e_{1}=1.$ In Eq. (2.59), the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
@@ 3.1 Fermionic bound states
From a spectral perspective, bound states correspond to a purely imaginary continuation of the quasi-momentum into the complex plane, $k\rightarrow i\kappa$, and therefore appear as isolated eigenvalues lying above and below the continuum threshold at $\pm M$. Accordingly, Eqs. (2.12) and (2.13) imply that the $u$ and $v$ components of a fermionic bound state decay as $e^{-\kappa|x|}$ as $|x|\rightarrow+\infty$.
@@ 3.2 Phase shifts and Levinson’s theorem
Figure 4: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $u(x)$ for $x<0$, $Re[c_{1}u_{in}(x)+c_{2}u_{ref}(x)]$ (green), and for $x>0$, $Im[u_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
@@
Figure 6: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $v(x)$ for $x<0$, $Re[c_{1}v_{in}(x)+c_{2}v_{ref}(x)]$ (green), and for $x>0$, $Re[v_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 7: (color online) Phase shift $\delta(k)$ vs $k$ of the scattering states $u$ and $v$ for the fermion mass $M=2.15\times 10^{-5}$. Note that $\delta(0)-\delta(+\infty)=\frac{\pi}{2}$.
@@
One therefore concludes that the upper and lower components of the spinor acquire identical phase shifts after scattering off the sine-Gordon–type soliton. This result contrasts with the recent findings in the ATM model with a soliton background of variable topological charge, where the phase shifts of the $u$ and $v$ components differ by a constant. In that case, a single phase shift is defined as the average of the phase shifts of the upper and lower components [2]. In the Fig. 5 we plot the common phase shift $\delta(k)=\delta_{u}(k)=\delta_{v}(k)$. Using this Fig. one can check the Levinson’s theorem [12]

### v2

@@ 3 Wronskian method and matching scattering states
The solutions to the differential equation (2.23) can be expressed in terms of the local Heun functions [4, 7, 5]. Note that we have used the parameter set provided in (B.6)-(B.8) for $e_{1}e_{2}=-1,e_{1}=1.$ In Eq. (2.76), the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
@@ 3.1 Fermionic bound states
The origin of these dispersion relations are discussed in the Appendix B.1, Eqs. (B.1)-(B.3). From a spectral perspective, bound states correspond to a purely imaginary continuation of the quasi-momentum into the complex plane, $k\rightarrow i\kappa$, and therefore appear as isolated eigenvalues lying above and below the continuum threshold at $\pm M$. Accordingly, Eqs. (2.22) and (2.23) imply that the $u$ and $v$ components of a fermionic bound state decay as $e^{-\kappa|x|}$ as $|x|\rightarrow+\infty$.
@@ 3.2 Phase shifts and Levinson’s theorem
Figure 4: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $u(x)$ for $x<0$, $Re[c_{1}u_{in}(x)+c_{2}u_{ref}(x)]$ (magenta), and for $x>0$, $Im[u_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
@@
Figure 6: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $v(x)$ for $x<0$, $Re[c_{1}v_{in}(x)+c_{2}v_{ref}(x)]$ (magenta), and for $x>0$, $Re[v_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 7: (color online) Phase shift $\delta(k)$ vs $k$ of the scattering states $u$ and $v$ for the fermion masses $M=2.15\times 10^{-5}$ (red), $5\times 10^{-5}$ (green), $10^{-4}$ (brown). Note that $\delta(0)-\delta(+\infty)=\frac{\pi}{2}$.
@@
One therefore concludes that the upper and lower components of the spinor acquire identical phase shifts after scattering off the sine-Gordon–type soliton. This result contrasts with the recent findings in the ATM model with a soliton background of variable topological charge, where the phase shifts of the $u$ and $v$ components differ by a constant. In that case, a single phase shift is defined as the average of the phase shifts of the upper and lower components [2]. In the Fig. 7 we plot the common phase shift $\delta(k)=\delta_{u}(k)=\delta_{v}(k)$ for three values of $M$. Using this Fig. one can check the Levinson’s theorem [12]
@@ 4 Discussions
Appendix A The sequence of transformations: $x\rightarrow y\rightarrow\theta\rightarrow w$
Next we provide the sequence of transformations $x\rightarrow y\rightarrow\theta\rightarrow w$ in order to get the Fuchsian second order differential equation (2.38) starting from the equation (2.23). The following identities can be written by applying successively the derivation chain-rule in each case.
1. The variable transformation $x\rightarrow y$: $y=-\tanh{(2Kx)}$ implies the next identities
$\frac{d}{dx}$ $=$ $-2K(1-y^{2})\frac{d}{dy},$ (A.1)
$\frac{d^{2}}{dx^{2}}$ $=$ $-8K^{2}y(1-y^{2})\frac{d}{dy}+4K^{2}(1-y^{2})^{2}\frac{d^{2}}{dy^{2}}.$ (A.2)
2. The variable transformation $y\rightarrow\theta$: $y=\cos{\theta}$ implies
$\frac{d}{dy}$ $=$ $-\frac{1}{\sin{\theta}}\frac{d}{d\theta},$ (A.3)
$\frac{d^{2}}{dy^{2}}$ $=$ $-\frac{\cos{\theta}}{\sin{\theta}^{3}}\frac{d}{d\theta}+\frac{1}{\sin{\theta}^{2}}\frac{d^{2}}{d\theta^{2}}.$ (A.4)
3. The variable transformation $\theta\rightarrow w$: $w=e^{i\theta}$ implies
$\frac{d}{d\theta}$ $=$ $iw\frac{d}{dw},$ (A.5)
$\frac{d^{2}}{d\theta^{2}}$ $=$ $-\frac{d}{dw}-w^{2}\frac{d^{2}}{dw^{2}}.$ (A.6)
Appendix B Relationships between the parameters: Cases 1 and 2
In this appendix we discuss the relationships between the relevant parameters.
B.1 Asymptotic scattering and bound states
For the scattering states one has
$E_{1}^{2}=k^{2}+M^{2}\rightarrow E_{1}=\pm\sqrt{k^{2}+M^{2}}.$ (B.1)
This is a dispersion relation which can be obtained from the asymptotic form of the equation (2.23) at $x\rightarrow\pm\infty$, i.e.
$u^{\prime\prime}+(E_{1}^{2}-M^{2})u=0.$ (B.2)
So, (B.1) follows for the free wave solution $u\sim e^{ikx}$ of (B.2). Similarly, for the spinor bound states one has
$E_{bs}^{2}=-\kappa^{2}+M^{2}\rightarrow E_{bs}=\pm\sqrt{M^{2}-\kappa^{2}}.$ (B.3)
It follows after substituting the expression of the form $u\sim e^{-\kappa x}\,(\kappa>0)$ for a wave with vanishing b.c. at $x\rightarrow+\infty$ into (B.2). The threshold states occur at $\kappa=0$ ($E_{thr}=\pm M$) and the zero modes at $\kappa=\pm M$ ($E_{bs}=0$).
B.2 Case 1
The variable transformation (2.40) $w\rightarrow z$: $z=\frac{w+1}{2}$ implies
$\frac{d}{dw}$ $=$ $\frac{1}{2}\frac{d}{dz},\,\,\,\,\frac{d^{2}}{dw^{2}}=\frac{1}{4}\frac{d^{2}}{dz^{2}}.$ (B.4)
Using these identities one can write the Fuchsian equation (2.38) in the form (2.42). This transformation sends $w=-1\rightarrow z=0,\,\,\,\,w=1\rightarrow z=1,\,\,\,\,w=0\rightarrow z=\frac{1}{2},\,\,\,\,w=\infty\rightarrow z=\infty$. So, the regular singularities on the complex plane $z$ become at
$z=0,1,a,\infty;\,\,\,\,\mbox{with}\,\,a=\frac{1}{2}.$ (B.5)
Next, we write the parameter values which have been considered in the work.
The canonical form of the Heun equation, given by (1.1)-(1.2), with the standard singular points in (B.5), is specified by the parameters $\alpha,\beta,\gamma,\delta,q,\epsilon$. These parameters are determined by the parameter sets
$\sigma$ $=$ $0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1,$ (B.6)
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=i\frac{E_{1}+e_{1}k}{K}$ (B.7)
$\gamma$ $=$ $\delta^{\star}=1-e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=-e_{1}\frac{ik}{2K},$ (B.8)
where $\delta^{\star}$ stands for complex conjugation of $\delta$.
B.3 Case 2
The variable transformation (2.77) $w\rightarrow z$: $z=\frac{1-w}{2}$ implies
$\frac{d}{dw}$ $=$ $-\frac{1}{2}\frac{d}{dz},\,\,\,\,\frac{d^{2}}{dw^{2}}=\frac{1}{4}\frac{d^{2}}{dz^{2}}.$ (B.9)
Using these identities one can write the Fuchsian equation (2.38) in the form (2.79). This transformation sends
$w=-1\rightarrow z=1,\,\,\,\,w=1\rightarrow z=0,\,\,\,\,w=0\rightarrow z=1/2,\,\,\,\,w=-\infty\rightarrow z=\infty.$ (B.10)
So, the regular singularities on the complex plane $z$ become the same as in (B.5). Similarly, in this case the canonical form of the Heun equation, given by (1.1)-(1.2), with the standard singular points in (B.5), is specified by the parameters $\alpha,\beta,\gamma,\delta,q,\epsilon$. These parameters are determined by the parameter sets
$\sigma$ $=$ $0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1.$ (B.11)
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=-i\frac{E_{1}+e_{1}k}{K}$ (B.12)
$\gamma$ $=$ $\delta^{\star}=1+e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=e_{1}\frac{ik}{2K}.$ (B.13)
Finally, observe that the parameter sets (B.6)-(B.8) corresponding to Case 1 and (B.11)-(B.13) corresponding to Case 2 are related through the transformations $E_{1}\leftrightarrow-E_{1}$ and $e_{1}\leftrightarrow-e_{1}$. This correspondence is consistent with the equivalence of the Fuchsian equations in Case 1, (2.42), and Case 2, (2.79), under the transformation $E_{1}\leftrightarrow-E_{1}$.